OpenStudy (anonymous):

Which is equivalent to (3 + 2i)(5 + i)?

OpenStudy (anonymous):

For this problem it looks like you have to foil the numbers. So you get: 13+13i

OpenStudy (lgbasallote):

"which" lol

OpenStudy (callisto):

(3 + 2i)(5 + i) = (3)(5 + i) + 2i (5 + i) = 15 + 3i + 10i + 2i^2 = 15 + 13i -2 = 13 + 13i

OpenStudy (anonymous):

13+13i's the answer. Any other issues?

OpenStudy (lgbasallote):

i do...i think it should be 13(i + 1) :P haha jk